A plane took 3h 33min to finish a journey.if the distance of the journey was 2840 miles at what average speed did the plane travel?
1
Expert's answer
2020-05-26T12:52:14-0400

The equation is easy, simply divide total distance by total time. The calculations are a bit more complex because we need to convert units. Most likely average speed in miles per second or minute doesn't tell us anything because we measure speeds in miles an hour.


We have 3 hours and 33 minutes. To convert 33 minutes to hours, divide 33 by 60 because 1 hour contains 60 minutes. Therefore, the average speed in milers per hour is


"v=\frac{d_\text{elapsed}}{t_\text{elapsed}}=\frac{2840\text{ miles}}{3\text{ h}+33\text{ min}/60\frac{\text{min}}{\text{h}}}=800\text{ mph}."